# Overhead in Quantum Circuits with Time-Multiplexed Qubit Control

### v1

@@ III.3 Scaling behavior
Here, we further explore the functional dependence of the serialization overhead on the dominant parameters. As could be expected, we observe that the overhead increases linearly with the number of single-qubit gates $N_{1}$ in the routed circuit, and the single-qubit gate time $t_{\text{1q}}$. We saw this behavior in Fig. 2(a) and Fig. 2(a) for $\times5$ and $\times11$ grids, respectively. We further show that this linear dependence holds for the $\times11$ grid with random circuits in Fig. 5 and with circuits from MQT Bench in Fig. 6, where data points are compared to the predicted linear function (solid lines) for different numbers of qubits per switch.
@@
Figure 7: Relative overhead by serialization compared to the routed circuit duration for different ratios of two-qubit to single-qubit gate durations, $t_{\text{2q}}$ and $t_{\text{1q}}$, respectively, for selected algorithms on an $\times11$ grid. (a) Shor’s algorithm. (b) QAOA. (c) Graph-state preparation. For increasing $t_{\text{2q}}$, the overhead by serialization decreases as serialized single-qubit gates are more likely to be concurrent with a two-qubit gate. The insets show the single- and two-qubit gate densities of the circuit. Note the logaritmic $x$ axis. Each marker represents the median over 30 seeds.
@@
While these observations explain a sublinear scaling in $k$, the logarithmic form is not immediately obvious. It can be rationalized using arguments from queueing theory, where extreme-value methods are applied to analyze maximum waiting times. In queueing models, the waiting time or queue length is treated as a random variable, and in many standard settings, the asymptotic behavior of the queue-length distribution has an exponentially decaying tail [82]. These methods are also used to analyze multiplexing and other types of signal transmission in classical communication networks [83, 84]. Practically, these properties of queues mean that the probability of encountering a very long queue decreases exponentially with its length, making long delays rare. In Appendix E, we present a heuristic derivation of how this leads to logarithmic scaling. Alternatively, in queueing-theory terms, each switch in our time-multiplexing setup can be viewed as a server that sequentially processes single-qubit gates (jobs) from multiple qubits (clients). The serialization of gates on a given switch generates waiting times, and the total circuit duration is determined by the maximum waiting time across all switches. It can be calculated that the expectation value of the maximum waiting time is logarithmic in the number of clients (qubits) [85].
@@
For the simulations above, we selected a ratio between single- and two-qubit gate durations $t_{\text{2q}}=10\,t_{\text{1q}}$ (see Sec. II.4.1). In Fig. 7, we present the relative overhead for different $t_{\text{2q}}/t_{\text{1q}}$ ratios for Shor’s algorithm, QAOA, and graph-state preparation. We see from the plots that larger $t_{\text{2q}}/t_{\text{1q}}$ ratios result in lower overhead, likely because a larger number of serialized single-qubit gates can be executed in parallel when there are long two-qubit gates to effectively hide the serialization cost. The functional scaling with qubits per switch $k$ does not appear to be affected by different ratios; it remains approximately logarithmic but with different multiplicative factors.
@@ IV Summary and conclusion
The trade-off we quantified occurs since a scalable architecture with shared control lines limits the possibility of gate parallelization. This leads to increased circuit runtimes, increasing exposure to decoherence, and the impact on algorithm execution therefore needs to be assessed. To this end, we quantified the serialization overhead for a range of standard benchmark quantum algorithms and random quantum circuits on three NISQ devices: a $\times5$-qubit square grid, an $\times11$-qubit square grid, and IBM’s 127-qubit Eagle heavy-hexagon layout. We assumed a hierarchy of time scales common to such devices: two-qubit gates take longer to execute than single-qubit gates, which in turn are slower than the switching delay in multiplexed control.
@@ V Outlook
Another natural direction for future work is to integrate serialization directly into routing and compilation algorithms, enabling global optimization of both qubit placement and control-line allocation. Here, factors such as coherence times varying across a quantum processor could be taken into account. The setup we studied could also be combined with frequency-multiplexing to further reduce the number of lines [90].
@@ Appendix A Compilation strategies for minimizing serialization overhead for single-qubit gates
Here, we provide further details about the algorithms we considered and used for minimizing serialization overhead for single-qubit gates. We can reduce the overhead from switch delays significantly by leveraging the temporal overlap between switching operations and concurrent two-qubit gates. Since two-qubit gates have durations equivalent to multiple single-qubit operations plus switching time in most situations we consider in this article (see Sec. II.4.1), the transition of control between qubits can often be scheduled during these longer operations without extending the total circuit duration. Our serialization model always places a zero-duration two-qubit switch gate between the source qubit and the target qubit to enforce temporal dependencies. When an actual two-qubit gate already exists between these qubits, the switch gate introduces no additional constraint. More importantly, if a two-qubit gate involves one qubit in the switching pair but not the other, the temporal dependency created by Sw does not cause serialization since coupler control is distinct from the qubit control, and the zero duration of the switch gate ensures no increase in total execution time.
The switch delay gate SDel would normally increase circuit duration by the switch response time. However, this delay can be hidden when switching occurs during concurrent two-qubit operations. We implement two optimization strategies to imitate this effect: (a) omit the delay gate when both qubits involved in the switch are simultaneously engaged in two-qubit gates, and (b) omit the delay gate when the temporal position of a neighboring two-qubit gate exceeds the temporal position of the single-qubit predecessor plus the switching duration. These optimizations ensure that control transitions are scheduled within existing idle periods, minimizing the impact on overall circuit execution time and representing a realistic switch scheduling.
@@ Appendix C Additional results for square-grid qubit layouts
Figure 11: Impact of time-multiplexing on execution of quantum algorithms from the MQT Bench set for a $\times5$ grid. (a) Total circuit duration for serialized routed circuits (colored markers) and routed circuits without serialization (black crosses and dashed line) across different quantum algorithms, ordered by increasing gate count. (b) Relative overhead of serialization compared to the routed circuit duration (markers, left axis), with shaded areas indicating gate densities ($\rho_{1}$ pink, $\rho_{2}$ blue, $\rho_{\rm total}$ green; right axis) and different marker colors indicating different numbers of qubits per switch. The data points are medians over 20 transpiler seeds. The full names of the quantum algorithms are listed in Table 2 in Appendix E.
@@
First, in Fig. 11, we show the MQT Bench results for the $\times5$ grid, in the same style as for the $\times11$ grid in Fig. 3 in the main text. Figure 11(a) shows the total circuit duration for different numbers of qubits per switch for the quantum algorithms in MQT Bench, while Fig. 11(b) shows the serialization overhead for the same cases, relative to the duration of the routed circuit. The results for the $\times5$ grid display the same patterns as those for the $\times11$ grid: serialization overhead varies widely between algorithms, but tends to be larger for algorithms with higher gate densities. However, we note that the absolute and relative serialization overhead is smaller for the $\times5$ grid than for the $\times11$ grid, even for the same number of qubits per switch. We attribute the difference in absolute serialization overhead to the fact that the benchmark circuits are longer for more qubits; for the difference in relative serialization overhead, we note that it is consistent with the pattern seen in the bottom row in Fig. 2, as discussed at the end of Sec. III.1.
Figure 12: Serialization overhead, in seconds, for the cases of 2 (blue squares) and 4 (orange circles) qubits per switch, for quantum algorithms from the MQT Bench set on a (a) $\times5$ and (b) $\times11$ square grid of qubits. Exact numerical values for $k=2$ are shown in Table 1. The full names of the quantum algorithms are listed in Table 2 in Appendix E.
Next, we zoom in on the results in Fig. 11(a) and plot the duration of only the serialization overhead for the cases of 2 and 4 qubits per switch in Fig. 12(a). We provide the same type of plot for the $\times11$ grid by zooming in on Fig. 3(a) in the main text to plot the serialization overhead for 2 and 4 qubits per switch in Fig. 12(b). These low numbers of qubits per switch are the most likely to be used in first implementations of time-multiplexed qubit control, which makes extracting the duration of the serialization overhead particularly important here, allowing comparison to state-of-the-art qubit coherence time to judge the feasibility of using switches.
@@
Figure 13: Distribution of time-multiplexing overhead across switch sizes on a $\times5$ grid, for the quantum algorithms from MQT Bench shown in Fig. 11. The histograms show the relative overhead (ratio of serialized to routed circuit duration) for different numbers of qubits per switch. The +n+n notation to the right in the bottom right histogram indicates $n$ additional counts beyond the histogram limits.
@@ Appendix D Results for a heavy-hexagon qubit layout
Figure 14: Overhead from time-multiplexing single-qubit control for a 127-qubit heavy-hexagon architecture. Each data point displayed is the median over 100 seeds creating random circuits; shaded areas show the interquartile range. (a) Serialization overhead as increase in circuit duration compared to the routed circuit, as a function of the number of gates in random circuits. (b) The same serialization overhead as in (a), but now shown relative to the overhead introduced by the qubit routing, as a function of the number of gates in random circuits. (c) Breakdown of total circuit duration: translated circuit (black), qubit routing overhead (light blue), and serialization overhead (blue, green, yellow), for three different numbers of gates in random circuits. The procedure for generating and compiling the circuits is detailed in Sec. II.
In Fig. 14, we show the overhead due to time-multiplexing for random circuits, in the way as we did for square grids in Fig. 2 in the main text. We observe the same patterns for the heavy-hexagon layout in Fig. 14 as for the square grids in Fig. 2: in Fig. 14(a), we see that the serialization overhead increases linearly with the number of gates in the circuits; in Fig. 14(b), we see that the ratio between the duration of the serialized circuits and the duration of the routed circuit remains constant as a function of the number of gates in the circuits (after some minor variation at the start); in Fig. 14(c), we see that routing itself gives a substantial overhead compared to the uncompiled circuit, but that having more than 2 qubits per switch gives an even larger overhead. Compared to the results for the $\times11$ square-grid layout in the right column of Fig. 2, the absolute serialization overhead is longer here for the 127-qubit heavy-hexagon layout, but the relative overhead compared to routing is quite similar. The reason for the increased routing overhead is the sparser connectivity of the heavy-hexagon layout compared to the square-grid layout.
Figure 15: Impact of time-multiplexing on execution of quantum algorithms from the MQT Bench set for a 127-qubit heavy-hexagon layout. (a) Total circuit duration for serialized routed circuits (colored markers) and routed circuits without serialization (black crosses and dashed line) across different quantum algorithms, ordered by increasing gate count. (b) Relative overhead of serialization compared to the routed circuit duration ((markers, left axis)), with shaded areas indicating gate densities ($\rho_{1}$ pink, $\rho_{2}$ blue, $\rho_{\rm total}$ green; right axis) and different marker colors indicating different numbers of qubits per switch. The data points are medians over 20 transpiler seeds. The full names of the quantum algorithms are listed in Table 2 in Appendix E.
@@
We now zoom in on the results in Fig. 15(a) and plot the duration of only the serialization overhead for the cases of 2 and 4 qubits per switch in Fig. 16, in the same way as we did for the $\times5$ grid in Fig. 12(a) and for the $\times11$ grid in Fig. 12(b). We see that the results for the 127-qubit heavy-hexagon layout are quite similar to those for the $\times11$ grid, with slightly longer serialization overheads in the heavy-hexagon case.
Figure 17: Distribution of time-multiplexing overhead across switch sizes on a 127-qubit heavy-hexagon architecture, for the quantum algorithms from MQT Bench shown in Fig. 15. The histograms show the relative overhead (ratio of serialized to routed circuit duration) for different numbers of qubits per switch. The +n+n notation to the right in the some histograms indicates $n$ additional counts beyond the histogram limits.

### v2

@@ III.3 Scaling behavior
We further explore the functional dependence of the serialization overhead on the dominant parameters. As could be expected, we observe that the overhead increases linearly with the number of single-qubit gates $N_{1}$ in the routed circuit, and the single-qubit gate time $t_{\text{1q}}$. We saw this behavior in Fig. 2(a) and Fig. 2(b) for $\times5$ and $\times11$ grids, respectively. We further show that this linear dependence holds for the $\times11$ grid with random circuits in Fig. 5 and with circuits from MQT Bench in Fig. 6, where data points are compared to the predicted linear function (solid lines) for different numbers of qubits per switch.
@@
Figure 7: Relative overhead by serialization compared to the routed circuit duration, as a function of qubits per switch, for different ratios of two-qubit to single-qubit gate durations, $t_{\text{2q}}$ and $t_{\text{1q}}$, respectively, for selected algorithms on an $\times11$ grid. (a) Shor’s algorithm. (b) QAOA. (c) Graph-state preparation. For increasing $t_{\text{2q}}$, the overhead by serialization decreases as serialized single-qubit gates are more likely to be concurrent with a two-qubit gate. The insets show the single- and two-qubit gate densities of the circuit. Note the logaritmic $x$ axis. Each marker represents the median over 30 seeds.
@@
While these observations explain a sublinear scaling in $k$, the logarithmic form is not immediately obvious. It can be rationalized using arguments from queueing theory, where extreme-value methods are applied to analyze maximum waiting times. In queueing models, the waiting time or queue length is treated as a random variable, and in many standard settings, the asymptotic behavior of the queue-length distribution has an exponentially decaying tail [92]. These methods are also used to analyze multiplexing and other types of signal transmission in classical communication networks [93, 94]. Practically, these properties of queues mean that the probability of encountering a very long queue decreases exponentially with its length, making long delays rare. In Appendix E, we present a heuristic derivation of how this leads to logarithmic scaling. Alternatively, in queueing-theory terms, each switch in our time multiplexing setup can be viewed as a server that sequentially processes single-qubit gates (jobs) from multiple qubits (clients). The serialization of gates on a given switch generates waiting times, and the total circuit duration is determined by the maximum waiting time across all switches. It can be calculated that the expectation value of the maximum waiting time is logarithmic in the number of clients (qubits) [95].
III.4 Effect of gate and switch durations
@@
For the simulations described above, we selected a ratio between single- and two-qubit gate durations $t_{\text{2q}}=10\,t_{\text{1q}}$ (see Sec. II.4.1). In Fig. 7, we present the relative overhead for different $t_{\text{2q}}/t_{\text{1q}}$ ratios for Shor’s algorithm, QAOA, and graph-state preparation. The plots show that larger $t_{\text{2q}}/t_{\text{1q}}$ ratios result in lower overhead, likely because more single-qubit gates can be executed serially during the longer two-qubit gates effectively hiding the serialization cost. The functional scaling with qubits per switch $k$ does not appear to be affected by different ratios; it remains approximately logarithmic, but with different multiplicative factors.
Figure 8: Relative overhead by serialization compared to the routed circuit duration, as a function of qubits per switch, for different switch delay times $t_{\text{sw}}$, for selected algorithms on an $\times11$ grid with $t_{\text{2q}}=10\,t_{\text{1q}}$. (a) Shor’s algorithm. (b) QAOA. (c) Graph-state preparation. Note the logarithmic $x$ axis. Each marker represents the median over 30 seeds and shaded areas correspond to the inter-quartile range. The insets show serialization overhead as a function of $t_{\text{sw}}$ for 2, 10 and 121 qubits per switch, color corresponding to the vertical dashed lines in the main plot. These plots confirms a linear dependence of the relative overhead on the switch delay time.
All circuits considered so far have used a switch delay time $t_{\text{sw}}=$2\text{\,}\mathrm{n}\mathrm{s}$\$, corresponding to the time it takes for a switching event to occur. This is fast but realistic for state-of-the-art devices [33]. The impact of longer switch times of up to $100\text{\,}\mathrm{n}\mathrm{s}$, modeled by the SDel gate in the serialized circuit, is shown in Fig. 8, again for Shor’s algorithm, QAOA, and graph-state preparation on the $\times11$ square grid. From Fig. 8, we see that the overhead depends linearly on the switch time, with steeper increase when there are more qubits per switch. The scaling of the overhead with the number of qubits per switch remains approximately logarithmic.
@@ IV Summary and conclusion
The trade-off we quantified occurs since a scalable architecture with shared control lines limits the possibility of gate parallelization. This leads to increased circuit runtimes, increasing exposure to decoherence, and therefore the impact on algorithm execution needs to be assessed. To this end, we quantified the serialization overhead for a range of standard benchmark quantum algorithms and random quantum circuits on three NISQ devices: a $\times5$-qubit square grid, an $\times11$-qubit square grid, and IBM’s 127-qubit Eagle heavy-hexagon layout. We assumed a hierarchy of time scales common to such devices: two-qubit gates take longer to execute than single-qubit gates, which in turn are slower than the switching delay in multiplexed control. We verified that our scaling results were robust to variations of the specific times when they respected this hierarchy.
@@ V Outlook
When scaling up with the goal of building a fault-tolerant quantum computer, the impact of time multiplexing on quantum error correction should be studied further using the framework we have introduced here. For a fixed number of control lines, there would be a trade-off between how large code distance one could achieve by using many qubits and how much the serialization overhead would degrade the performance of such a longer-distance code. The analysis of such a trade-off would depend heavily on details of the hardware, e.g., gate times for different types of gates, the error budget for these gates, and other error sources. For particular hardware layouts and error-correction codes, there can be highly optimized arrangements of operations in time and space [100].
Another natural direction for future work is to integrate serialization directly into routing and compilation algorithms, enabling global optimization of both qubit placement and control-line allocation. Here, factors such as coherence times varying across a quantum processor could be taken into account. The setup we studied could also be combined with frequency multiplexing to further reduce the number of lines [101].
@@ Acknowledgements.
Data Availability Statement
The code and data that support the findings of this article will be openly available at https://github.com/marvin-richter/overhead-time-multiplexing upon publication.
@@ Appendix A Compilation strategies for minimizing serialization overhead for single-qubit gates
Here, we provide further details about the algorithms we considered and used for minimizing serialization overhead for single-qubit gates. We can reduce the overhead from switch delays significantly by leveraging the temporal overlap between switching operations and concurrent two-qubit gates. Since two-qubit gates have durations equivalent to multiple single-qubit operations plus switching time in most situations we consider in this article (see Sec. II.4.1), the transition of control between qubits can often be scheduled during these longer operations without extending the total circuit duration. Our serialization model always places a zero-duration two-qubit switch gate between the source qubit and the target qubit to enforce temporal dependencies. When an actual two-qubit gate already exists between these qubits, the switch gate does not introduce an additional constraint. More importantly, if a two-qubit gate involves one qubit in the switching pair but not the other, the temporal dependency created by Sw does not cause serialization since the coupler control is distinct from the qubit control, and the zero duration of the switch gate ensures that there is no increase in total execution time.
The switch delay gate SDel would normally increase the circuit duration by the switch response time. However, this delay can be hidden when switching occurs during concurrent two-qubit operations. We implement two optimization strategies to imitate this effect: (a) omit the delay gate when both qubits involved in the switch are simultaneously engaged in two-qubit gates, and (b) omit the delay gate when the temporal position of a neighboring two-qubit gate exceeds the temporal position of the single-qubit predecessor plus the switching duration. These optimizations ensure that control transitions are scheduled within existing idle periods, minimizing the impact on overall circuit execution time and representing a realistic switch scheduling.
@@ Appendix C Additional results for square-grid qubit layouts
Figure 12: Impact of time multiplexing on execution of quantum algorithms from the MQT Bench set for a $\times5$ grid. (a) Total circuit duration for serialized routed circuits (colored markers) and routed circuits without serialization (black crosses and dashed line) across different quantum algorithms, ordered by increasing gate count. (b) Relative overhead of serialization compared to the routed circuit duration (markers, left axis), with shaded areas indicating gate densities ($\rho_{1}$ pink, $\rho_{2}$ blue, $\rho_{\rm total}$ green; right axis) and different marker colors indicating different numbers of qubits per switch. The data points are medians over 20 transpiler seeds. The full names of the quantum algorithms are listed in Table 2 in Appendix E.
@@
First, in Fig. 12, we show the MQT Bench results for the $\times5$ grid, in the same style as for the $\times11$ grid in Fig. 3 in the main text. Figure 12(a) shows the total circuit duration for different numbers of qubits per switch for the quantum algorithms in MQT Bench. Fig. 12(b) shows the relative increase in circuit duration due to serialization (the ratio of serialized to routed circuit duration). The results for the $\times5$ grid display the same patterns as those for the $\times11$ grid: serialization overhead varies widely between algorithms, but tends to be larger for algorithms with higher gate densities. However, we note that the absolute and relative serialization overhead is smaller for the $\times5$ grid than for the $\times11$ grid, even for the same number of qubits per switch. We attribute the difference in absolute serialization overhead to the fact that the benchmark circuits are longer for more qubits; for the difference in relative serialization overhead, we note that it is consistent with the pattern seen in the bottom row in Fig. 2, as discussed at the end of Sec. III.1.
Figure 13: Serialization overhead, in seconds, for the cases of 2 (blue squares) and 4 (orange circles) qubits per switch, for quantum algorithms from the MQT Bench set on a (a) $\times11$ and (b) $\times5$ square grid of qubits. Exact numerical values for $k=2$ are shown in Table 1. The full names of the quantum algorithms are listed in Table 2 in Appendix E.
Next, we zoom in on the results in Fig. 3(a) in the main text and plot the duration of only the serialization overhead for the cases of 2 and 4 qubits per switch in Fig. 13(a). We provide the same type of plot for the $\times5$ grid by zooming in on Fig. 12(a) to plot the serialization overhead for 2 and 4 qubits per switch in Fig. 13(b). These low numbers of qubits per switch are the most likely to be used in first implementations of time-multiplexed qubit control, which makes extracting the duration of the serialization overhead particularly important here, allowing comparison to state-of-the-art qubit coherence time to judge the feasibility of using switches.
@@
Figure 14: Distribution of time multiplexing overhead across switch sizes on a $\times5$ grid, for the quantum algorithms from MQT Bench shown in Fig. 12. The histograms show the relative overhead (ratio of serialized to routed circuit duration) for different numbers of qubits per switch. The +n+n notation to the right in the bottom right histogram indicates $n$ additional counts beyond the histogram limits.
@@ Appendix D Results for a heavy-hexagon qubit layout
Figure 15: Overhead from time multiplexing single-qubit control for a 127-qubit heavy-hexagon architecture. Each data point displayed is the median over 100 seeds creating random circuits; shaded areas show the interquartile range. (a) Serialization overhead as increase in circuit duration compared to the routed circuit, as a function of the number of gates in random circuits. (b) The same serialization overhead as in (a), but now shown relative to the duration of the routed circuit, as a function of the number of gates in random circuits. (c) Breakdown of total circuit duration: translated circuit (black), qubit routing overhead (light blue), and serialization overhead (blue, green, yellow), for three different numbers of gates in random circuits. The procedure for generating and compiling the circuits is detailed in Sec. II.
In Fig. 15, we show the overhead due to time multiplexing for random circuits, in the way as we did for square grids in Fig. 2 in the main text. We observe the same patterns for the heavy-hexagon layout in Fig. 15 as for the square grids in Fig. 2: in Fig. 15(a), we see that the serialization overhead increases linearly with the number of gates in the circuits; in Fig. 15(b), we see that the ratio between the duration of the serialized circuits and the duration of the routed circuit remains constant as a function of the number of gates in the circuits (after some minor variation at the start); in Fig. 15(c), we see that routing itself gives a substantial overhead compared to the uncompiled circuit, but that having more than 2 qubits per switch gives an even larger overhead. Compared to the results for the $\times11$ square-grid layout in the right column of Fig. 2, the absolute serialization overhead is longer here for the 127-qubit heavy-hexagon layout, but the relative overhead compared to routing is quite similar. The reason for the increased routing overhead is the sparser connectivity of the heavy-hexagon layout compared to the square-grid layout.
Figure 16: Impact of time multiplexing on execution of quantum algorithms from the MQT Bench set for a 127-qubit heavy-hexagon layout. (a) Total circuit duration for serialized routed circuits (colored markers) and routed circuits without serialization (black crosses and dashed line) across different quantum algorithms, ordered by increasing gate count. (b) Relative overhead of serialization compared to the routed circuit duration ((markers, left axis)), with shaded areas indicating gate densities ($\rho_{1}$ pink, $\rho_{2}$ blue, $\rho_{\rm total}$ green; right axis) and different marker colors indicating different numbers of qubits per switch. The data points are medians over 20 transpiler seeds. The full names of the quantum algorithms are listed in Table 2 in Appendix E.
@@
We now zoom in on the results in Fig. 16(a) and plot the duration of only the serialization overhead for the cases of 2 and 4 qubits per switch in Fig. 17, in the same way as we did for the $\times11$ grid in Fig. 13(a) and for the $\times5$ grid in Fig. 13(b). We see that the results for the 127-qubit heavy-hexagon layout are quite similar to those for the $\times11$ grid, with slightly longer serialization overheads in the heavy-hexagon case.
Figure 18: Distribution of time multiplexing overhead across switch sizes on a 127-qubit heavy-hexagon architecture, for the quantum algorithms from MQT Bench shown in Fig. 16. The histograms show the relative overhead (ratio of serialized to routed circuit duration) for different numbers of qubits per switch. The +n+n notation to the right in the some histograms indicates $n$ additional counts beyond the histogram limits.